OpenStudy (anonymous):

I am trying to simplify this : 2(x-1)(x+1)^2+2(x+1)(x-1)^2

myininaya (myininaya):

did you ever look at my attachment to your earlier post?

jhonyy9 (jhonyy9):

(x+1)(x-1)(2(x+1)+2(x-1))=(x+1)(x-1)(2(x+1+x-1)=(x+1)(x-1)(2(2x))= =(x+1)(x-1)4x